At a price of $2.50 per bushel, the annual US supply and demand for corn are 8.5 and 9.8 million bushels,respectively. When the price rises to $3.30, the supply increases to 10.5 million bushels, while the demand decreases to 7.8 bushels.
A) assuming that the price-supply and the price-demand equations are linear, find equations for each.
B) Find the equilibrium point for the US corn market.

Answer by ewatrrr(24785)

A)($3.30, 10.5)
($2.50, 8.5) m = 2/.8 = 2.5
y[s] - 8.5 = 2.5(x -2.5)
y[s] = 2.5x + 8.5 -6.25
.......
($2.50, 10.5)
(3.50, 7.8) m = 2.7/-1 = -2.7
y[d] -10.5 = -2.7(x - 2.5)
y[d] = -2.7x + 10.5 + 6.75
B) y[s] = y[d]
2.5x + 2.25 = -2.7x + 17.25
x = 15/5.2 = 2.8846 million bushels